# PlanetServe: A Decentralized, Scalable, and Privacy-Preserving Overlay for Democratizing Large Language Model Serving

### v4

@@ Abstract
While significant progress has been made in research and development on open-source and cost-efficient large-language models (LLMs), serving scalability remains a critical challenge, particularly for small organizations and individuals seeking to deploy and test their LLM innovations. Inspired by peer-to-peer networks that leverage decentralized overlay nodes to increase throughput and availability, we propose PlanetServe, an LLM serving overlay that harnesses computing resources from decentralized contributors. We identify four key research problems inherent to enabling such a decentralized infrastructure: 1) overlay network organization; 2) LLM communication privacy; 3) overlay forwarding for resource efficiency; and 4) verification of serving quality. This work presents the first systematic study of these fundamental problems in the context of decentralized LLM serving. Evaluation results from a prototype implemented on a set of decentralized nodes demonstrate that PlanetServe achieves a latency reduction of over 50% compared to the baseline design without overlay forwarding. Furthermore, the security features introduce minimal overhead to serving latency and throughput. We believe this work pioneers a new direction for democratizing and scaling future AI serving capabilities.
@@ 1 Introduction
The rapid development and proliferation of open-source, cost-efficient large-language models (LLMs)–such as Google’s Gemma [gemma], DeepSeek-R1 [deepseek], and Meta’s LLaMA series [touvron2023llama]–have significantly lowered the barrier for deploying AI systems. This progress makes it feasible to run advanced language models on personal workstations and local servers with consumer GPUs. Despite this accessibility, serving scalability remains a critical challenge for small organizations and individuals to deploy their LLM innovations. Individual researchers, university laboratories, and start-up companies often find it prohibitively expensive and technically challenging to deploy their models at scale, particularly when popularity surges quickly. For instance, some promising models initially provided excellent performance at low cost, but later suffered severe performance degradation when user demand suddenly increased [DeepseekSlow]. Traditional cloud-based infrastructures, which are not always affordable or accessible for smaller organizations and individuals.
We propose PlanetServe, a decentralized overlay infrastructure designed for large-scale LLM deployment. Inspired by PlanetLab [planetlab], a well-known peer-to-peer (P2P) network allowing researchers to deploy and test Internet applications across distributed nodes, PlanetServe is designed to harness computing resources from a diverse range of contributors, including individuals, academic institutions, and small organizations. Participants in PlanetServe volunteer their computing power and network resources, and in return gain the ability to deploy their own LLMs and AI-driven services on this shared, decentralized network.
@@
PlanetServe thus represents a novel approach to deploying LLMs and other advanced AI services, democratizing large-scale AI capabilities and enhancing user privacy. Our position is that model serving today is mostly concentrated in large-scale cloud deployments, but the software ecosystem [mlc-llm] has already demonstrated that inference can be effectively deployed across a wide range of devices. This trend suggests that the future inference landscape should also be determined by cost efficiency, privacy, and availability. A decentralized infrastructure like PlanetServe is well-positioned to adapt to such a landscape: whether the dominant models are large or small, participants can contribute heterogeneous resources that align with their needs (e.g., low-latency private inference or specialized model serving).
@@
We implemented PlanetServe’s core functions as a prototype consisting of decentralized nodes and tested the prototype in a public cloud without central control. This approach provided a controllable testbed, necessary as we do not have direct access to a large set of physical decentralized servers across the Internet. We let each node add synthetic latency to every packet for the wide-area Internet conditions. Upon acceptance of this paper, we will make the prototype implementation and datasets open-sourced. We plan to deploy and maintain PlanetServe at scale in the future.
This work focuses exclusively on the LLM serving problem in a decentralized overlay network and assumes model training is completed by model developers. We do not consider decentralized LLM training [ye2024openfedllm] or collaborative serving an LLM instance [borzunov2023petal, DistServe2024, SplitWise2024]. This work is also different from distributed LLM serving in a cloud with centralized scheduling [sglang24, Preble25, DistServe2024, SplitWise2024, CacheGen2024].
The rest of this paper is organized as follows. Section 2 presents the system models and design objectives of PlanetServe. Section 3 details the design of PlanetServe. We provide the security analysis and evaluation in Section 4. The performance evaluation results are presented in Section 5. We discuss the related work in Section 7. We conclude this work and discuss possible future work in Section 8.
@@ 2.2 Cost Advantage and Incentive Model
Cost advantage of PlanetServe. The cost advantage of PlanetServe compared to commercial public cloud is mainly from two facts. First, computation resources remain dramatically underutilized in organizations. Hence, the cost of contributing servers to PlanetServe is very low. Many organizations deployed GPU servers for ML training, but the servers sit idle outside of training times. It has been reported that the average server utilization rate is between 12%-18% of capacity [underulti]. 10 million servers are estimated to sit completely idle, representing \$30 billion in wasted capital [underulti]. These idle servers can be contributed to PlanetServe and form a large-scale serving infrastructure. Second, using community-contributed GPUs for inference can dramatically cut costs compared to commercial cloud platforms. Measurements show that the energy cost of running LLM inference on GPUs such as the V100 and A100 is relatively modest [samsi2023words], implying that the high hourly rates of public cloud instances are driven largely by infrastructure and pricing overheads rather than intrinsic hardware or energy expenses. Recent work on scalable inference similarly reports $20\times$ lower cost when serving LLMs on consumer GPUs versus AWS instances [ScaleLLM]. In addition, many consumer GPUs or GPUs that have been given away from training can still serve moderately sized LLMs (7B–13B parameters).
In summary, prior measurements and cost analyses suggest that LLM serving and testing using private hardware can be an order of magnitude cheaper than compared to public cloud if the hardware is fully utilized. However, scaling private hardware is challenging, and hardware utilization is very low. PlanetServe addresses both the scalability and under-utilization problem of private hardware by allowing different organizations to contribute their resources and share them with each other.
@@ 2.3 Threat Model and Assumptions
  • A compromised verification node could provide wrong information to users or false verification results for model nodes. We assume an adversary, potentially leveraging Sybil attacks to control multiple verification nodes. However, the adversary cannot control more than compromise more than $f$ out of $N=3f+1$ total verification nodes following the Byzantine fault tolerance (BFT) assumptions. This threshold is based on the assumption that verification nodes are run by decentralized entities, making the cost of compromising a sufficient fraction extremely high.
@@ 3.1 Overview
The user nodes form a dynamic overlay network. Upon joining PlanetServe, a user node registers its current IP address and public key with the verification nodes, whose IP addresses and public keys are public information. The public key serves as the identifier. It then downloads a list of active user and model nodes from a verification node, and the list is signed by more than 2/3 of the verification nodes [PBFT99]. Verification nodes may choose to divide the whole system into multiple regions and create a list of users and model nodes for each region, only when the number of users in each region is sufficiently large to hide the requester’s identity, for example, >1000>1000 users [Intruder, hordes00]. PlanetServe allows a user to use other users as relays for anonymous communication: each user request will be forwarded along an overlay path to hide the address of the original sending user, and the LLM reply will also be forwarded along the anonymous path back to the requesting user.
@@
The committee of verification nodes uses a recent BFT consensus protocol [TendermintBFT] to ensure consistency. The committee maintains a list of user nodes registered to PlanetServe, including their public keys, and a list of model nodes in PlanetServe, including their public keys and reputation scores. The reputation score of a model node quantifies the quality of its LLM responses and will be defined in Section 3.4. Verification nodes periodically request the model nodes anonymously and update their scores based on the responses.
@@ 3.2 Anonymous Overlay and Content Privacy
We analyze typical overlay anonymous communication protocols [Tor04, IS-NSDI07, LAP12, garlic], including Onion routing and sliced routing [IS-NSDI07, garlic], in the Appendix and find that none of them fit the above requirements. If we apply Onion to the overlay network in PlanetServe, there are a few main challenges. 1) The first layer of Onion nodes (called guards in Tor) will definitely know the sending user of a request, which should be avoided in PlanetServe. 2) The failure probability of an Onion path increases exponentially with the path length, especially for an overlay network of users with a high churn rate. 3) Each relay needs to perform public-key decryption on the entire message, introducing non-trivial overhead for users.
@@ 3.3 Overlay Forwarding among Model Nodes
State synchronization. For each model node in a group, it periodically broadcasts the local updates of its HR-tree, Each node keeps a recent snapshot of its HR-tree and the following updates after the snapshot. The node periodically sends a minimal but necessary update to all nodes in the domain.
Load balancing is another design consideration in addition to KV cache reusing. The main objectives of load balancing are to ensure that no model node receives more requests than its capacity, and every model node can serve a certain number of requests to establish a reputation score. In the forwarding logic shown in Fig.4, when selecting the target model nodes from a set of candidates, we always select the one with the lowest load balance (LB) factor, which is defined as $F_{LB}=L*(Q/C)$, where $L$ is the moving average of the latency to serve each request on that node, $Q$ is the number of queued requests on that node that are waiting for serving, $C$ is the capacity of the number of concurrent requests that can be served on the node. The moving average is calculated similarly to the RTT estimation with parameter $\alpha=1/8$. The LB factor is calculated locally by each node and audited by the verification committee. Each node periodically broadcasts its LB factor to other model nodes.
@@
Session affinity for consecutive prompts. Every model node includes its IP address in the response message. For subsequent follow-up prompts from the same user session, the user attempts to direct these prompts specifically to that designated model node, routing them via the anonymous overlay. This policy prioritizes KV cache reuse for consecutive prompts within a session.
@@ 4.1 Anonymity
We apply a commonly used entropy-based metric [zhuang2005cashmere, garlic] to measure the anonymity of a system, whose definition is presented in the Appendix. It ranges from 0 and 1, and a higher value represents better anonymity.
@@ 4.4 Other Security Properties
Denial-of-Service (DoS) attacks. We consider two scenarios of DoS attacks. 1) A malicious verification node refuses to perform the verification task, either acting as the leader for an epoch or as a member. If the leader is malicious and refuses to send challenge prompts to model nodes or broadcasts challenge responses to the committee, the system simply fails the current epoch and restores the verification process in the next epoch with a new leader. If some committee members refuse to verify and vote, the system can still rely on the remaining honest verification nodes to vote and reach a consensus, which is guaranteed by BFT. 2) A model node does not respond to user requests. This behavior can be detected by the verification process.
@@
Sybil attacks. An adversary could potentially leverage Sybil attacks to control multiple verification nodes to alter verification results. Verification nodes are operated by organizations or individuals with verifiable credentials (e.g., recognized academic or research institutions). These credentials will be periodically checked and revoked if a verification node is compromised, and then the node will be expelled from the committee. Moreover, verification nodes are selected to be geographically and administratively diverse, which makes it hard for an adversary to compromise $\geq 1/3$ of the committee. To further improve the security of the verification committee, the committee members can be non-static. They are periodically rotated and re-elected. Misbehaving verification nodes will be detected and removed. This could increase the cost and difficulty for the adversary to compromise the committee.
@@ 5.1 Methodology
We deploy the testbed of PlanetServe in a public cloud. The testbed includes eight model nodes with mid-tier hardware, each carrying one NVIDIA RTX A6000 GPU (48 GB) and serving a Meta-Llama-3 8B model, eight other model nodes with high-performance hardware, each carrying one NVIDIA A100 GPU (80 GB) and serving a DeepSeek-R1-Qwen-14B model. We run two machines as the verification nodes. One carries an NVIDIA A100 (40 GB SXM4), 30 virtual CPUs, 200 GiB of RAM, and a 0.5 TiB SSD, and the other carries an NVIDIA GH200 (96 GB HBM), 64 virtual CPUs, 432 GiB of RAM, and 4 TiB SSD. There is no centralized coordinator in the testbed. Internet latency in simulations is from samples collected by our measurement. We use the most up-to-date version of vLLM [vllm, vllmgithub], a recently developed LLM serving library, on each single model node.
@@
1. ToolUse (ToolBench[guo2024stabletoolbench]): .210k queries with tool-specific instructions; prompts average 7,206 tokens after Zipf-1.1 sampling. Prefix sharing is moderate, and outputs are capped at 100 tokens per query for the models we evaluate.
2. Coding (APPS[hendrycksapps2021]): .10,000 coding problems with detailed solution requests; prompts average 1,802 tokens after Zipf-0.8 sampling. Prefix overlap is minimal, and outputs are capped at 1,000 tokens per query for the models we evaluate.
3. Long-Doc QA (LooGLE[li2023loogle]): 776 long documents paired with 6.4k questions. Each prompt includes a long document as prefix followed by a question, averaging 10,985 tokens per prompt after Zipf-0.6 sampling. Outputs are capped at 100 tokens per query for the models we evaluate.
@@
Upon acceptance of this paper, we will make the prototype and data sets open-sourced.
@@ 5.2 Anonymous Routing
We evaluate the performance overhead of anonymous message (clove) processing using the Toolbench dataset [guo2024stabletoolbench]. Clove preparation is performed on a model node equipped with a single NVIDIA A100 40GB SXM4 GPU, 30 vCPUs, 200 GiB RAM, and 0.5 TiB SSD running Llama-3.3-70B. Decryption is executed on a user node with a single Intel i7-7700 CPU @ 3.60GHz. Figure 13 shows the cumulative distribution functions (CDFs) for both preparation and decryption latency over 10,000 trials.
Clove preparation is computationally lightweight, with a mean latency of 0.273 ms and a median (P50) latency of 0.280 ms. Even at the P99 percentile, latency remains below 0.31 ms, demonstrating tight performance bounds. Clove decryption is similarly efficient, with a slightly higher mean latency of 0.302 ms and a P50 of 0.309 ms. The P99 latency is 0.334 ms, and the success rate remains 100%, confirming the reliability of the decryption process. Overall, both operations exhibit low and stable latency.
We conduct simulations to evaluate the communication survival rate under network churn in a 3,119-node network and a churn rate of 200 nodes/min – a very high rate. We deployed 14 AWS t3.micro instances across multiple regions and measured end-to-end routing latency with realistic prompt payloads over 4,000 runs (3 nodes randomly fail in the USA; 1 randomly fails worldwide). We incorporate latency, link failures, packet loss, and congestion to the simulation and show the results in Fig. 13. The path survival rate of PlanetServe is the highest among these methods. When there is no failure, all methods have high delivery rates. However, when failure occurs, only PlanetServe can maintain a similar high delivery rate. The delivery rate of Onion significantly decreases.

### v5

@@ Abstract
While significant progress has been made in research and development on open-source and cost-efficient large language models (LLMs), serving scalability remains a critical challenge, particularly for small organizations and individuals seeking to deploy and test their LLM innovations. Inspired by peer-to-peer networks that leverage decentralized overlay nodes to increase throughput and availability, we propose PlanetServe, an LLM serving overlay that harnesses computing resources from decentralized contributors. We identify four key research problems inherent to enabling such a decentralized infrastructure: 1) overlay network organization; 2) LLM communication privacy; 3) overlay forwarding for resource efficiency; and 4) verification of serving quality. This work presents the first systematic study of these fundamental problems in the context of decentralized LLM serving. Evaluation results from a prototype implemented on a set of decentralized nodes demonstrate that PlanetServe achieves a latency reduction of over 50% compared to the baseline design without overlay forwarding. Furthermore, the security features introduce minimal overhead to serving latency and throughput. We believe this work pioneers a new direction for democratizing and scaling future AI serving capabilities.
@@ 1 Introduction
The rapid development and proliferation of open-source, cost-efficient large-language models (LLMs), such as Google’s Gemma [gemma], DeepSeek-R1 [deepseek], and Meta’s LLaMA series [touvron2023llama], have significantly lowered the barrier to AI deployment. This progress makes it feasible to run advanced language models on personal workstations and local servers with consumer GPUs. Despite this accessibility, serving scalability remains a critical challenge for small organizations and individuals to deploy LLM innovations. Individual researchers, university laboratories, and start-up companies often find it prohibitively expensive and technically challenging to deploy models at scale, particularly when popularity surges. For instance, some promising models initially provided excellent performance at low cost, but later suffered severe performance degradation when user demand suddenly increased [DeepseekSlow]. Traditional cloud-based infrastructures are not always affordable or accessible for smaller organizations and individuals.
We propose PlanetServe, a decentralized overlay infrastructure designed for large-scale LLM deployment. Inspired by PlanetLab [planetlab], a well-known peer-to-peer (P2P) network allowing researchers to deploy and test Internet applications across distributed nodes, PlanetServe harnesses computing resources from a diverse range of contributors, including individuals, academic institutions, and small organizations. Participants in PlanetServe volunteer computing power and network resources, and in return gain the ability to deploy LLMs and AI-driven services on this shared, decentralized network.
@@
PlanetServe represents a novel approach to deploying LLMs and other advanced AI services, democratizing large-scale AI capabilities while enhancing user privacy. Although model serving today is largely concentrated in large-scale cloud deployments, the software ecosystem [mlc-llm] has demonstrated that inference can be effectively deployed across diverse devices. This trend suggests a future inference landscape should also be determined by cost efficiency, privacy, and availability. A decentralized infrastructure like PlanetServe is well-suited to this evolution, enabling participants to contribute heterogeneous resources align with their needs (e.g., low-latency private inference or specialized model serving).
@@
We implemented PlanetServe’s core functions as a prototype consisting of decentralized nodes and tested the prototype in a public cloud without central control. This approach provided a controllable testbed, necessary as we do not have direct access to a large set of physical decentralized servers across the Internet. We let each node add synthetic latency to every packet for the wide-area Internet conditions. Upon acceptance of this paper, we will make the prototype implementation and datasets open-sourced.
This work focuses on decentralized LLM serving over an overlay network and assumes model training is completed by developers. We exclude decentralized training [ye2024openfedllm], collaborative serving of a LLM instance [borzunov2023petal, DistServe2024, SplitWise2024], and cloud-based distributed serving with centralized scheduling [sglang24, Preble25, DistServe2024, SplitWise2024, CacheGen2024].
The rest of this paper is organized as follows. Sec. 2 presents the system model and design objectives of PlanetServe. Sec. 3 details the system design. Sec. 4 and Sec. 5 provide the security analysis and performance evaluation, respectively. Sec. 6 discusses design insights and practical considerations, followed by Sec. 7, which outlines the limitations of the current approach. Sec. 8 reviews related work. Finally, Sec. 9 concludes the paper and discusses future directions.
@@ 2.2 Cost Advantage and Incentive Model
Cost advantage of PlanetServe. The cost advantage of PlanetServe compared to commercial public cloud is mainly from two facts. First, computation resources remain dramatically underutilized in organizations. Hence, the cost of contributing servers to PlanetServe is very low. Many organizations deployed GPU servers for ML training, but the servers sit idle outside of training times. It has been reported that the average server utilization rate is between 12%-18% of capacity [underulti]. 10 million servers are estimated to sit completely idle, representing \$30 billion in wasted capital [underulti]. These idle servers can be contributed to PlanetServe and form a large-scale serving infrastructure. Second, using community-contributed GPUs for inference can dramatically cut costs compared to commercial cloud platforms. Measurements show that the energy cost of running LLM inference on GPUs such as the V100 and A100 is relatively modest [samsi2023words], implying that the high hourly rates of public cloud instances are driven largely by infrastructure and pricing overheads rather than intrinsic hardware or energy expenses. Recent work on scalable inference similarly reports $20\times$ lower cost when serving LLMs on consumer GPUs versus AWS instances [ScaleLLM]. In addition, many consumer GPUs or GPUs that have been given away from training can still serve moderately sized LLMs (7B–13B parameters). Notably, this comparison focuses on cost efficiency, whereas commercial public clouds provide additional services beyond cost.
In summary, prior measurements and cost analyses suggest that LLM serving and testing using private hardware can be an order of magnitude cheaper than public cloud if the hardware is fully utilized. However, scaling private hardware is challenging, and hardware utilization is very low. PlanetServe addresses both the scalability and under-utilization problem of private hardware by allowing different organizations to contribute their resources and share them with each other.
@@ 2.3 Threat Model and Assumptions
  • A compromised verification node could provide wrong information to users or false verification results for model nodes. We assume an adversary, potentially leveraging Sybil attacks to control multiple verification nodes. No more than $f$ out of $N=3f+1$ verification nodes can be compromised following the Byzantine fault tolerance (BFT) assumption. This threshold is based on the fact that verification nodes are run by decentralized entities, making the cost of compromising a sufficient fraction extremely high.
@@ 3.1 Overview
The user nodes form a dynamic overlay network. Upon joining PlanetServe, a user node registers its current IP address and public key with the verification nodes, whose IP addresses and public keys are public information. The public key serves as the identifier. It then downloads a list of active user and model nodes from a verification node, and is signed by more than 2/3 of the verification nodes [PBFT99]. Verification nodes may choose to divide the whole system into multiple regions and create a list of users and model nodes for each region, only when the number of users in each region is sufficiently large to hide the requester’s identity, for example, >1000>1000 users [Intruder, hordes00]. PlanetServe allows a user to use other users as relays for anonymous communication: each user request is forwarded along an overlay path to hide the address of the original sender, and the LLM reply will also be forwarded along the anonymous path back to the requesting user.
@@
The committee of verification nodes uses a recent BFT consensus protocol [TendermintBFT] to ensure consistency. The committee maintains a list of user nodes registered to PlanetServe, including their public keys, and a list of model nodes in PlanetServe, including their public keys and reputation scores. The reputation score of a model node quantifies the quality of its LLM responses and will be defined in Sec. 3.4. Verification nodes periodically request the model nodes anonymously and update their scores based on the responses.
@@ 3.2 Anonymous Overlay and Content Privacy
We analyze typical anonymous overlay communication protocols [Tor04, Han2006RR, IS-NSDI07, LAP12, garlic], including Onion routing and sliced routing [Han2006RR, IS-NSDI07, garlic], in the Appendix and find that none of them fit the above requirements. If we apply Onion to the overlay network in PlanetServe, there are a few main challenges. 1) The first layer of Onion nodes (called guards in Tor) will definitely know the sending user of a request, which should be avoided in PlanetServe. 2) The failure probability of an Onion path increases exponentially with the path length, especially for an overlay network of users with a high churn rate. 3) Each relay needs to perform public-key decryption on the entire message, introducing non-trivial overhead for users.
@@ 3.3 Overlay Forwarding among Model Nodes
State synchronization. For each model node in a group, it periodically broadcasts the local updates of its HR-tree, each node keeps a snapshot of its HR-tree and the following updates after the snapshot. The node periodically sends a minimal but necessary update to all nodes in the group. Temporary inconsistencies (e.g., due to churn) during synchronization may reduce cache hit rates without affecting correctness since routing is constrained to nodes running the requested model.
Load balancing is another design consideration alongside KV cache reuse. Its main objectives are to prevent overload and ensure each model serves enough requests to establish a reputation score. As shown in Fig.4, target nodes are selected by minimizing the load-balance (LB) factor, $F_{LB}=L*(Q/C)$, where $L$ is the moving average of service latency, $Q$ the queued request number, and CC is the capacity of the number of concurrent requests. The moving average follows RTT estimation with $\alpha=1/8$. LB factors are computed locally, audited by the verification committee, and periodically broadcast to other nodes. Consequently, request routing based on $F_{LB}$ adapts to runtime variations across model nodes. As a node’s hardware performance degrades or if a node is inherently slower (e.g., consumer-grade GPUs), its $L$ value increases, which redirects incoming requests to nodes with higher effective capacity.
@@
Session affinity for consecutive prompts. Every model node includes its IP address in the response message. For subsequent prompts from the same user session, the user routes these prompts specifically to that designated model node, routing them via the anonymous overlay. This policy prioritizes KV cache reuse for consecutive prompts within a session.
@@ 4.1 Anonymity
We apply a commonly used entropy-based metric [zhuang2005cashmere, garlic] to measure the anonymity of a system, whose definition is presented in the Appendix A5. Definition of Anonymity Metric. It ranges from 0 and 1, and a higher value represents better anonymity.
@@ 4.4 Other Security Properties
Denial-of-Service (DoS) attacks. We consider three DoS scenarios. 1) If the leader is malicious and refuses to send challenge prompts to model nodes or broadcasts challenge responses to the committee, the system simply fails the current epoch and restores the verification process in the next epoch with a new leader. 2) If some committee members refuse to verify and vote, the system can still rely on the remaining honest verification nodes to vote and reach a BFT consensus. 3) A model node does not respond to user requests. This behavior can be detected by the verification protocol.
@@
Sybil attacks. An adversary could potentially leverage Sybil attacks to control multiple verification nodes to alter verification results. PlanetServe requires verification nodes to pass credential checks and jointly validate committee decisions via BFT consensus, eliminating trust on a single authority. Verification nodes are operated by entities with real-world identities (e.g., recognized academic or research institutions), which imposes non-trivial cost on large-scale Sybil creation. Compromised nodes can be revoked and removed upon detection. Moreover, verification nodes are selected to be geographically and administratively diverse, making it difficult for an adversary to compromise $\geq 1/3$ of the committee. To further limit prolonged adversarial influence, committee members are periodically rotated through randomized re-selection, and misbehaving nodes are excluded.
@@ 5.1 Methodology
We deploy the testbed of PlanetServe in a public cloud. The testbed includes eight model nodes with mid-tier hardware, each carrying one NVIDIA RTX A6000 GPU (48 GB) and serving a Meta-Llama-3 8B model, eight other model nodes with high-performance hardware, each carrying one NVIDIA A100 GPU (80 GB) and serving a DeepSeek-R1-Qwen-14B model. We run two machines as the verification nodes. One carries an NVIDIA A100 (40 GB SXM4), 30 virtual CPUs, 200 GiB of RAM, and a 0.5 TiB SSD, and the other carries an NVIDIA GH200 (96 GB HBM), 64 virtual CPUs, 432 GiB of RAM, and 4 TiB SSD. There is no centralized coordinator. Internet latency in simulations is drawn from our measurements. We use the latest version of vLLM [vllm, vllmgithub], a recently developed LLM serving library, on each single model node.
@@
1. ToolUse (ToolBench[guo2024stabletoolbench]): 210k queries with tool-specific instructions; prompts average 7,206 tokens after Zipf-1.1 sampling. Prefix sharing is moderate, and outputs are capped at 100 tokens per query for the models we evaluate.
2. Coding (APPS[hendrycksapps2021]): 10,000 coding problems with detailed solution requests; prompts average 1,802 tokens after Zipf-0.8 sampling. Prefix overlap is minimal, and outputs are capped at 1,000 tokens per query for the models we evaluate.
3. Long-Doc QA (LooGLE[li2023loogle]): 776 long documents paired with 6.4k questions. Each prompt includes a long document as a prefix followed by a question, averaging 10,985 tokens per prompt after Zipf-0.6 sampling. Outputs are capped at 100 tokens per query for the models we evaluate.
@@ 5.2 Anonymous Routing
We evaluate the performance overhead of anonymous message (clove) processing using the Toolbench dataset [guo2024stabletoolbench]. Clove preparation is performed on a model node equipped with a single NVIDIA A100 40GB SXM4 GPU, 30 vCPUs, 200 GiB RAM, and 0.5 TiB SSD running Llama-3.3-70B. Decryption is executed on a user node with a single Intel i7-7700 CPU @ 3.60GHz. Figure 13 shows the CDFs for both preparation and decryption latency over 10,000 trials.
Clove preparation is lightweight, with mean and P50 latencies of 0.273 and 0.280 ms. Even at the P99 percentile, latency remains below 0.31 ms, indicating tight bounds. Clove decryption is similarly efficient, with a slightly higher mean latency of 0.302 ms and a P50 of 0.309 ms. The P99 latency is 0.334 ms, while maintaining a 100% success rate. Overall, both operations exhibit low and stable latency.
We simulate communication survival under churn in a 3,119-node network and a churn rate of 200 nodes/min – a very high rate and deployed 14 AWS t3.micro instances across multiple regions and measured end-to-end routing latency with realistic prompt payloads over 4,000 runs (3 nodes randomly fail in the USA; 1 randomly fails worldwide). The simulation incorporates latency, link failures, packet loss, and congestion (Fig. 13). PlanetServe achieves the highest path survival rate, maintaining high delivery under failures, while Onion degrades significantly.